# D-TPT: Dimensional Entropy Maximization for Calibrating Test-Time Prompt Tuning in Vision-Language Models

### v1

@@ 1 Introduction
Furthermore, Test-time Prompt Tuning (TPT) extends its applicability to unlabeled settings by adapting prompts online (Shu et al., 2022). Although TPT effectively improves predictive accuracy, it simultaneously induces overconfidence due to its reliance on prediction entropy minimization as the objective function. Recent studies (Zhang et al., 2024b; Han et al., 2025) attempt to mitigate the degradation in model calibration performance caused by entropy minimization and improve practical applicability. Notably, calibration methods in test-time adaptation for VLMs are based on empirical observations. While increasing text feature diversity effectively improves calibration (Yoon et al., 2024; Sharifdeen et al., 2025); however diversity-based approaches suffers from a lack of understanding regarding its underlying mechanism.
@@
The correlation between average text feature dispersion (ATFD) and expected calibration error (ECE) (Naeini et al., 2015) has been demonstrated through various experiments (Yoon et al., 2024). Nevertheless, Figure 1 shows that calibration capability can be achieved despite low ATFD, suggesting that underlying causal factors remain to be uncovered. In particular, we conjecture that the modality gap (Liang et al., 2022) between image and text features is crucial role to this phenomenon. A detailed discussion of our geometric intuition is provided in Section 4.4.
Contrastive VLMs such as CLIP are trained by maximizing the cosine similarity between image and text embeddings. Nevertheless, a significant modality gap between image and text features is widely observed, and various approaches have been proposed to understand and exploit this phenomenon (Ouali et al., 2023; Oh et al., 2023). The modality gap is not expressed uniformly across all feature dimensions; instead, it is concentrated in a single dominant dimension. As depicted in Figure 2, modality gap arises from distinct basis in the image and text representations. These dominant dimensions are consistently observed across all text classes and all image samples. We denote the dimension prevailing in textual embeddings as the text-dominant dimension (TDD) and the corresponding one in visual embeddings as the image-dominant dimension (IDD).
@@ 2 Related Work
Prompt tuning. Pretrained models are utilized in various applications, and prompt tuning is a parameter-efficient fine-tuning approach that adapts large pretrained models to downstream tasks by training only a small number of input tokens (Lester et al., 2021). Prompt tuning has become a standard solution alongside LoRA (Hu et al., 2022) and adapters (Rebuffi et al., 2017; Zhang et al., 2021), demonstrating its effectiveness across vision models (Wang et al., 2022c; Jia et al., 2022; Bahng et al., 2022), and language models (Li & Liang, 2021; Liu et al., 2021). In VLMs, CoOp (Zhou et al., 2022b) demonstrates that input contexts such as “a photo of a [class]” significantly influence downstream performance. To overcome the limitations of manual prompt engineering, CoOp introduced context optimization, where prompts are trained rather than hand-crafted. Building on this idea, prompt tuning for VLMs has since been widely explored. Notably, CoCoOp (Zhou et al., 2022a) utilizes image features to generate image-conditioned prompts that enhance generalization to unseen classes, while MaPLe (Khattak et al., 2023) extends context tuning by applying prompts to both the image encoder and the text encoder.
@@
Test-time adaptation. Deep neural networks show strong performance in visual tasks but remain vulnerable under distribution shifts (Shimodaira, 2000). Test-time adaptation has been proposed to address dynamically changing real-world environments by enabling pretrained models to adapt during inference in an unsupervised manner, thereby avoiding the need for additional pretraining. Research on test-time adaptation has primarily focused on entropy minimization (Wang et al., 2021; Niu et al., 2023; Lee et al., 2024; Han et al., 2025) and consistency regularization (Wang et al., 2022a; Döbler et al., 2023; Liu et al., 2024). Under this paradigm, TPT (Shu et al., 2022) adopts entropy minimization through augmented input views and prompt tuning, yielding effective adaptation of VLMs during inference and providing a foundation for subsequent research. (Feng et al., 2023; Zhang et al., 2024a; Xiao et al., 2025; Sheng et al., 2025).
@@
Model calibration. Modern neural networks often suffer from overconfidence, which leads to miscalibration (Guo et al., 2017). Calibration denotes the alignment between predicted probabilities and the true likelihood of correctness, and it is essential for reliable deployment in real-world settings. Existing methods include post-hoc based approaches (Platt et al., 1999; Guo et al., 2017; Zhang et al., 2020) such as temperature scaling and regularization-based approaches (Kumar et al., 2018; Mukhoti et al., 2020) that refine the objective function. Given the calibration problems of test-time adaptation in VLMs, C-TPT (Yoon et al., 2024) improves diversity by encouraging text embeddings to disperse from the centroid. In contrast, SaLS (Zhang et al., 2024a) argues that miscalibration originates from the logit range and proposes a method that adjusts logits within the zero-shot range. Furthermore, O-TPT (Sharifdeen et al., 2025) ensures orthogonality across text features, which enhances class separability and mitigates directional bias.
@@ 3 Preliminary
Zero-shot classification. CLIP (Radford et al., 2021) is pretrained on massive image and text pairs to maximize the cosine similarity between the outputs of the image encoder $E_{visual}$ and the text encoder $E_{text}$. For zero-shot inference, the inputs to the image encoder and the text encoder are an image $x$ and a set of context prompts $\{[\mathrm{p};y_{c}]\}_{c=1}^{C}$, where $\mathrm{p}$ denotes the prompt and $y_{c}$ denotes the class context, with $C$ corresponds to the number of classes. Following previous work (Sharifdeen et al., 2025), we adopt the prompt “a photo of a [class]” as the base prompt. Accordingly, the predicted probability of class $c$ for an image $x_{i}$ is given by $p_{i,c}=\frac{\exp(\tau\cdot cos(t_{c},v_{i}))}{\sum_{k=1}^{C}\exp(\tau\cdot cos(t_{k},v_{i}))},$ where $v_{i}=E_{visual}(x_{i})$ is the image feature and $t_{c}=E_{text}([\mathrm{p};y_{c}])$ is the text feature, and $\tau$ denotes the learnable logit scaling factor.
@@
Test-time prompt tuning protocol. Following test-time prompt tuning approaches (Shu et al., 2022; Yoon et al., 2024; Sharifdeen et al., 2025), we aim to improve performance by updating the prompt $\mathrm{p}$ during inference on a single test sample $\{x_{i}^{test},y_{i}^{test})\in\{\mathcal{X}_{test},\mathcal{Y}_{test}\}$. Here, $\mathcal{Y}_{test}$ is only used for performance evaluation and is not accessed during test-time. During inference, the prompt is tuned in an unsupervised online manner. Specifically, TPT (Shu et al., 2022) generates additional N1N-1 augmented views for each test image, and optimizes the prompt by minimizing the entropy of high-confidence samples. The objective function is defined as
@@
where $H_{i}$ denotes the entropy (i.e., $H_{i}=-p_{i}\log p_{i}$), $\rho$ is a confidence threshold, and $N$ is the number of inputs. Algorithm 1 illustrates the overall procedure of test-time prompt tuning. Based on the initial zero-shot predictions, TPT minimizes the entropy of high-confidence samples, while our proposed D-TPT further improves calibration through the regularization loss function described in Section 4. The prompt is updated via gradient descent on the objective function. After adaptation and prediction for a single test sample, the model is reset to its initial weights.
@@ 4.1 Dimensional sensitivity
where $q=\sigma(\tau\cdot cos(t,v))$ denotes the original prediction probability, $\sigma(\cdot)$ denotes a softmax function, and $p_{m}$ denotes the prediction obtained by masking the $m\text{-}th$ dimension, $KLD(\cdot\mid\mid\cdot)$ denotes the Kullback–Leibler divergence. Figure 3(a) presents the sensitivity analysis across feature dimensions, which shows that dominant dimensions exhibit substantially higher sensitivity. Moreover, Figure 3(b) illustrates performance changes when dominant dimensions are replaced with their mean. For zero-shot CLIP, constraining these dimensions improves accuracy and calibration capability in some cases. For TDD, consistent reductions in average ECE were observed in both zero-shot CLIP and TPT. These results indicate that dominant dimensions have a pronounced impact on prediction, while their sensitivity can increase predictive uncertainty.
@@
4.2 D-TPT: Dimensional entropy maximiztion for test-time prompt tuning
@@ 4.4 Geometric support
where $z_{i,c}$ and $p_{i,c}$ denote logit and probability distribution corresponding to class $c$ for input $x_{i}$, respectively, and $\theta$ denotes the model parameters.
@@
Our proposed D-TPT reduces the influence of dominant basis in the image features, thereby narrowing the modality gap. By maximizing entropy across all feature dimensions, D-TPT amplifies the binormal directions directions on the hypersphere. Empirically, our method exhibits similar ATFD to TPT but with a lower logit range, suggesting diversity in binormal directions. Since the logit mean is proportional to the average cosine similarity between image and text feature, we can confirm the reduction of the modality gap. Therefore, the reduced logit range of D-TPT leads to a decrease in softmax probabilities, which mitigates overconfidence (Murugesan et al., 2024).
@@ 5.1 Experimental setup
Datasets. Following the standard evaluation protocol for downstream tasks in VLMs (Zhou et al., 2022b; Yoon et al., 2024), we use datasets under natural domain shift, including ImageNet-A (Hendrycks et al., 2021b), ImageNet-V2 (Recht et al., 2019), ImageNet-R (Hendrycks et al., 2021a), and ImageNet-Sketch (Wang et al., 2019). For fine-grained classification, we use ImageNet (Deng et al., 2009), DTD (Cimpoi et al., 2014), Flowers102 (Nilsback & Zisserman, 2008), Food101 (Bossard et al., 2014), SUN397 (Xiao et al., 2010), FGVC-Aircraft (Maji et al., 2013), Oxford Pets (Parkhi et al., 2012), Caltech101 (Li et al., 2022), UCF101 (Soomro et al., 2012), EuroSAT (Helber et al., 2018), and Stanford Cars (Krause et al., 2013). Experimental results are reported as the average performance across all datasets for both the natural domain shift and fine-grained classification datasets. Moreover, we present the mean and standard deviation over three trials with randoms 0, 1, and 2.
@@ 5.2 Experimental results
Fine-grained classification. Table 1 shows the average performance across eleven fine-grained datasets for evaluation on downstream tasks. Compared with zero-shot CLIP, TPT consistently shows improved accuracy but worse calibration performance. Designed for improved calibration, C-TPT, O-TPT, and D-TPT apply additional regularization for calibration. These calibration-oriented methods achieve improved accuracy than zero-shot CLIP while reducing calibration error compare to TPT. Among them, D-TPT achieves 0.74%, 0.60%, and $5.58\times 10^{3}$ improvements in accuracy, ECE, and AURC, respectively, compared to the previous sota O-TPT on CLIP-ViT-B/16. On the other hand, O-TPT demonstrates the best performance in MCE, indicating the need for calibration evaluation across the various metrics. We do not achieve the best performance on CLIP-RN50, however, we achieve competitive performance between C-TPT and O-TPT. Detailed results are provided in A.1.
Natural distribution shifts. Table 2 shows the average performance across four ImageNet variants, consisting ImageNet-A/R/V2/Sketch for out-of-distribution evaluation. For CLIP-ViT-B/16, we achieve 0.61% lower performance and 1.31% improved calibration error compared to the previous state-of-the-art O-TPT. Compared to the zero-shot CLIP, we achieve improved accuracy and ECE by 0.68% and 1.10%, respectively, demonstrating the potential for improvement across all metrics through test-time prompt tuning. Furthermore, for CLIP-RN50, it demonstrates improved performance across all metrics compared to existing methods. Detailed results are provided in A.2.
@@ 5.3 Further analysis
Failure case analysis. Figure 6 presents confidence histograms (above) and reliability diagrams (below) to analyze the distribution of predictions and calibration trends with respect to confidence. Following C-TPT (Yoon et al., 2024), we report average accuracy and confidence over 20 bins. From both success and failure cases in comparison with existing methods, we observe that D-TPT more closely preserves the confidence distribution of the zero-shot CLIP. Therefore, the beneficial effects of D-TPT are particularly pronounced in cases where TPT significantly increases overconfidence. Furthermore, this observation is supported by the geometric interpretation discussed in Section 4.4. Specifically, D-TPT aims to align all class text features with the image feature, hence mitigating changes in the features of non-top-1 classes induced by TPT. Whereas C-TPT alleviates such changes primarily for classes that are closer to the centroid than the top-1 prediction. Consequently, our method can be understood as particularly effective in scenarios where TPT amplifies calibration error, since it counteracts the update directions for a broader set of class text features.
@@
Pareto front analysis. Regularization methods for calibration improve calibration capability but often suffer accuracy degradation (Kumar et al., 2018; Karandikar et al., 2021; Yoon et al., 2024). Regarding the trade-off between accuracy and calibration performance, Figure 7 presents the variations in performance of C-TPT, O-TPT, and D-TPT according to the coefficient of the regularization term. Increasing marker size indicates a larger influence of the regularization loss with increasing hyperparameter $\lambda$. Experimental results demonstrate the superiority of the proposed method for a range of $\lambda$ values.

### v2

@@ 1 Introduction
Furthermore, Test-time Prompt Tuning (TPT) extends its applicability to unlabeled settings by adapting prompts online (Shu et al., 2022). Although TPT effectively improves predictive accuracy, it simultaneously induces overconfidence due to its reliance on prediction entropy minimization as the objective function. Recent studies (Zhang et al., 2024b; Han et al., 2025) attempt to mitigate the degradation in model calibration performance caused by entropy minimization and improve practical applicability. Notably, calibration methods in test-time adaptation for VLMs are based on empirical observations. While increasing text feature diversity effectively improves calibration (Yoon et al., 2024; Sharifdeen et al., 2025); however diversity-based approaches suffer from a lack of understanding regarding its underlying mechanism.
@@
The correlation between average text feature dispersion (ATFD) and expected calibration error (ECE) (Naeini et al., 2015) has been demonstrated through various experiments (Yoon et al., 2024). Nevertheless, Figure 1 shows that calibration capability can be achieved despite low ATFD, suggesting that underlying causal factors remain to be uncovered. In particular, we conjecture that the modality gap (Liang et al., 2022) between image and text features plays a crucial role in this phenomenon. A detailed discussion of our geometric intuition is provided in Section 4.4.
Contrastive VLMs such as CLIP are trained by maximizing the cosine similarity between image and text embeddings. Nevertheless, a significant modality gap between image and text features is widely observed, and various approaches have been proposed to understand and exploit this phenomenon (Ouali et al., 2023; Oh et al., 2023). The modality gap is not expressed uniformly across all feature dimensions; instead, it is concentrated in a single dominant dimension. As depicted in Figure 2, modality gap arises from distinct bases in the image and text representations. These dominant dimensions are consistently observed across all text classes and all image samples. We denote the dimension prevailing in textual embeddings as the text-dominant dimension (TDD) and the corresponding one in visual embeddings as the image-dominant dimension (IDD).
@@ 2 Related Work
Prompt tuning. Pretrained models are utilized in various applications, and prompt tuning is a parameter-efficient fine-tuning approach that adapts large pretrained models to downstream tasks by training only a small number of input tokens (Lester et al., 2021). Prompt tuning has become a standard solution alongside LoRA (Hu et al., 2022) and adapters (Rebuffi et al., 2017; Zhang et al., 2021), demonstrating its effectiveness across vision models (Wang et al., 2022c; Jia et al., 2022; Bahng et al., 2022) and language models (Li & Liang, 2021; Liu et al., 2021). In VLMs, CoOp (Zhou et al., 2022b) demonstrates that input contexts such as “a photo of a [class]” significantly influence downstream performance. To overcome the limitations of manual prompt engineering, CoOp introduced context optimization, where prompts are trained rather than hand-crafted. Building on this idea, prompt tuning for VLMs has since been widely explored. Notably, CoCoOp (Zhou et al., 2022a) utilizes image features to generate image-conditioned prompts that enhance generalization to unseen classes, while MaPLe (Khattak et al., 2023) extends context tuning by applying prompts to both the image encoder and the text encoder.
@@
Test-time adaptation. Deep neural networks show strong performance in visual tasks but remain vulnerable under distribution shifts (Shimodaira, 2000). Test-time adaptation has been proposed to address dynamically changing real-world environments by enabling pretrained models to adapt during inference in an unsupervised manner, thereby avoiding the need for additional pretraining. Research on test-time adaptation has primarily focused on entropy minimization (Wang et al., 2021; Niu et al., 2023; Lee et al., 2024; Han et al., 2025) and consistency regularization (Wang et al., 2022a; Döbler et al., 2023; Liu et al., 2024). Under this paradigm, TPT (Shu et al., 2022) adopts entropy minimization through augmented input views and prompt tuning, yielding effective adaptation of VLMs during inference and providing a foundation for subsequent research (Feng et al., 2023; Zhang et al., 2024a; Xiao et al., 2025; Sheng et al., 2025).
@@
Model calibration. Modern neural networks often suffer from overconfidence, which leads to miscalibration (Guo et al., 2017). Calibration denotes the alignment between predicted probabilities and the true likelihood of correctness, and it is essential for reliable deployment in real-world settings. Existing methods include post-hoc calibration approaches (Platt et al., 1999; Guo et al., 2017; Zhang et al., 2020) such as temperature scaling and regularization-based approaches (Kumar et al., 2018; Mukhoti et al., 2020) that refine the objective function. Given the calibration problems of test-time adaptation in VLMs, C-TPT (Yoon et al., 2024) improves diversity by encouraging text embeddings to disperse from the centroid. In contrast, SaLS (Murugesan et al., 2024) argues that miscalibration originates from the logit range and proposes a method that adjusts logits within the zero-shot range. Furthermore, O-TPT (Sharifdeen et al., 2025) promotes orthogonality across text features, which enhances class separability and mitigates directional bias.
@@ 3 Preliminary
Zero-shot classification. CLIP (Radford et al., 2021) is pretrained on massive image and text pairs to maximize the cosine similarity between the outputs of the image encoder $E_{visual}$ and the text encoder $E_{text}$. For zero-shot inference, the inputs to the image encoder and the text encoder are an image $x$ and a set of context prompts $\{[\mathrm{p};y_{c}]\}_{c=1}^{C}$, where $\mathrm{p}$ denotes the prompt and $y_{c}$ denotes the class context, where $C$ corresponds to the number of classes. Following previous work (Sharifdeen et al., 2025), we adopt the prompt “a photo of a [class]” as the base prompt. Accordingly, the predicted probability of class $c$ for an image $x_{i}$ is given by $p_{i,c}=\frac{\exp(\tau\cdot\cos(t_{c},v_{i}))}{\sum_{k=1}^{C}\exp(\tau\cdot\cos(t_{k},v_{i}))},$ where $v_{i}=E_{visual}(x_{i})$ is the image feature and $t_{c}=E_{text}([\mathrm{p};y_{c}])$ is the text feature, and $\tau$ denotes the learnable logit scaling factor.
@@
Test-time prompt tuning protocol. Following test-time prompt tuning approaches (Shu et al., 2022; Yoon et al., 2024; Sharifdeen et al., 2025), we aim to improve performance by updating the prompt $\mathrm{p}$ during inference on a single test sample $\{x_{i}^{test},y_{i}^{test}\}\in\{\mathcal{X}_{test},\mathcal{Y}_{test}\}$. Here, $\mathcal{Y}_{test}$ is only used for performance evaluation and is not accessed during test-time. During inference, the prompt is tuned in an unsupervised online manner. Specifically, TPT (Shu et al., 2022) generates additional N1N-1 augmented views for each test image, and optimizes the prompt by minimizing the entropy of high-confidence samples. The objective function is defined as
@@
where $H$ denotes the entropy (i.e., $H(p_{i})=-\sum_{c}p_{i,c}\log p_{i,c}$) is a confidence threshold, and $N$ is the number of inputs. Algorithm 1 illustrates the overall procedure of test-time prompt tuning. Based on the initial zero-shot predictions, TPT minimizes the entropy of high-confidence samples, while our proposed D-TPT further improves calibration through the regularization loss function described in Section 4. The prompt is updated via gradient descent on the objective function. After adaptation and prediction for a single test sample, the model is reset to its initial weights.
@@ 4.1 Dimensional sensitivity
where $q=\sigma(\tau\cdot\cos(t,v))$ denotes the original prediction probability, $\sigma(\cdot)$ denotes a softmax function, and $p_{m}$ denotes the prediction obtained by masking the $m\text{-}th$ dimension, $KLD(\cdot\mid\mid\cdot)$ denotes the Kullback–Leibler divergence. Figure 3(a) presents the sensitivity analysis across feature dimensions, which shows that dominant dimensions exhibit substantially higher sensitivity. Moreover, Figure 3(b) illustrates performance changes when dominant dimensions are replaced with their mean. For zero-shot CLIP, constraining these dimensions improves accuracy and calibration capability in some cases. For TDD, consistent reductions in average ECE were observed in both zero-shot CLIP and TPT. These results indicate that dominant dimensions have a pronounced impact on prediction, while their sensitivity can increase predictive uncertainty.
@@
4.2 D-TPT: Dimensional entropy maximization for test-time prompt tuning
@@ 4.4 Geometric support
where $z_{i,c}$ and $p_{i,c}$ denote logit and probability corresponding to class $c$ for input $x_{i}$, respectively, and $\theta$ denotes the model parameters.
@@
Our proposed D-TPT reduces the influence of dominant basis in the text features, thereby narrowing the modality gap. By maximizing entropy across all feature dimensions, D-TPT amplifies the binormal directions on the hypersphere. Empirically, our method exhibits similar ATFD to TPT but with a lower logit range, suggesting diversity in binormal directions. Since the logit mean is proportional to the average cosine similarity between image and text feature, we can confirm the reduction of the modality gap. Therefore, the reduced logit range of D-TPT leads to a decrease in softmax probabilities, which mitigates overconfidence (Murugesan et al., 2024).
@@ 5.1 Experimental setup
Datasets. Following the standard evaluation protocol for downstream tasks in VLMs (Zhou et al., 2022b; Yoon et al., 2024), we use datasets under natural domain shift, including ImageNet-A (Hendrycks et al., 2021b), ImageNet-V2 (Recht et al., 2019), ImageNet-R (Hendrycks et al., 2021a), and ImageNet-Sketch (Wang et al., 2019). For fine-grained classification, we use ImageNet (Deng et al., 2009), DTD (Cimpoi et al., 2014), Flowers102 (Nilsback & Zisserman, 2008), Food101 (Bossard et al., 2014), SUN397 (Xiao et al., 2010), FGVC-Aircraft (Maji et al., 2013), Oxford Pets (Parkhi et al., 2012), Caltech101 (Li et al., 2022), UCF101 (Soomro et al., 2012), EuroSAT (Helber et al., 2018), and Stanford Cars (Krause et al., 2013). Experimental results are reported as the average performance across all datasets for both the natural domain shift and fine-grained classification datasets. Moreover, we present the mean and standard deviation over three trials with random seeds 0, 1, and 2.
@@ 5.2 Experimental results
Fine-grained classification. Table 1 shows the average performance across eleven fine-grained datasets for evaluation on downstream tasks. Compared with zero-shot CLIP, TPT consistently shows improved accuracy but worse calibration performance. Designed for improved calibration, C-TPT, O-TPT, and D-TPT apply additional regularization for calibration. These calibration-oriented methods achieve improved accuracy than zero-shot CLIP while reducing calibration error compared to TPT. Among them, D-TPT achieves 0.74%, 0.60%, and $5.58\times 10^{3}$ improvements in accuracy, ECE, and AURC, respectively, compared to the previous state-of-the-art O-TPT on CLIP-ViT-B/16. On the other hand, O-TPT demonstrates the best performance in MCE, indicating the need for calibration evaluation across the various metrics. We do not achieve the best performance on CLIP-RN50, however, we achieve competitive performance between C-TPT and O-TPT. Detailed results are provided in A.1.
Natural distribution shifts. Table 2 shows the average performance across four ImageNet variants, consisting of ImageNet-A/R/V2/Sketch for out-of-distribution evaluation. For CLIP-ViT-B/16, we achieve 0.61% lower performance and 1.31% improved calibration error compared to the previous state-of-the-art O-TPT. Compared to the zero-shot CLIP, we achieve improved accuracy and ECE by 0.68% and 1.10%, respectively, demonstrating the potential for improvement across all metrics through test-time prompt tuning. Furthermore, for CLIP-RN50, D-TPT demonstrates improved performance across all metrics compared to existing methods. Detailed results are provided in A.2.
@@ 5.3 Further analysis
Failure case analysis. Figure 6 presents confidence histograms (above) and reliability diagrams (below) to analyze the distribution of predictions and calibration trends with respect to confidence. Following C-TPT (Yoon et al., 2024), we report average accuracy and confidence over 20 bins. From both success and failure cases in comparison with existing methods, we observe that D-TPT more closely preserves the confidence distribution of the zero-shot CLIP. Therefore, the beneficial effects of D-TPT are particularly pronounced in cases where TPT significantly increases overconfidence. Furthermore, this observation is supported by the geometric interpretation discussed in Section 4.4. Specifically, D-TPT aims to align all class text features with the image feature, hence mitigating changes in the features of non-top-1 classes induced by TPT. In contrast, C-TPT alleviates such changes primarily for classes that are closer to the centroid than the top-1 prediction. Consequently, our method can be understood as particularly effective in scenarios where TPT amplifies calibration error, since it counteracts the update directions for a broader set of class text features.
@@
Pareto front analysis. Regularization methods for calibration improve calibration capability but often suffer from accuracy degradation (Kumar et al., 2018; Karandikar et al., 2021; Yoon et al., 2024). Regarding the trade-off between accuracy and calibration performance, Figure 7 presents the variations in performance of C-TPT, O-TPT, and D-TPT according to the coefficient of the regularization term. Increasing marker size indicates a larger influence of the regularization term with increasing hyperparameter $\lambda$. Experimental results demonstrate the superiority of the proposed method for a range of $\lambda$ values.